Operating Systems and Systems Integration

Lab Topics for Part Time Course

1. Partition hard disk, install Linux Source: lab sheets on partitioning and installing Red Hat 7.3 from the network.

2. Introduction to the shell, Basic Tools Modules 1, 2 and 3 of workshop notes.

3. More Tools Modules 3 and 4 of workshop notes.

4. Shell Progamming 1 My lab sheets on shell programming.

5. Shell Progamming 2 My lab sheets on shell programming.

6. Secure Shell My chapter on Secure Shell in workshop notes.

7. Apache 1 My chapters on Apache in workshop notes.

8. Apache 2 My chapters on Apache in workshop notes.

9. NFS, Samba 1 My chapters on NFS in workshop notes. My worksheets on using Samba as a primary domain controller.

10. Samba 2 My worksheets on using Samba as a primary domain controller.

Nick Urbanik <nicku@vtc.edu.hk>  ver. 1.2